AAVSOnet status, 2016-02-07

Sun, 02/07/2016 - 18:53

SRO is down for a while.  A cable snagged and both rotated the camera as well as pulling a connector off of the focuser motor.  John is ordering a new motor, but there will be no new data after 2016-02-06 (and that data is out of focus).

OC61 is up, but lousy weather.  Recently they are having internet problems as well, so it is difficult to load new objects into the queue.  We are still waiting for the focal reducer for the spectrograph.

TMO61 remains down until the dome shutter is repaired.

coker30 is up, but lousy weather.  Should be improving for a few nights, though!

BSM_Hamren is up, but the queues need to be freshened to make better use of its time.

BSM_HQ is up; Helmar Adler is currently upgrading it to ACP V8/Expert as a test bed for our other systems.  I shoveled snow off of the enclosure yesterday, so we may get some new data soon if the weather permits.

BSM_NM is down while the mount is being repaired.

BSM_South is operational with a loaner ST10 camera from Peter Nelson, while the main ST8 camera is being repaired.

BSM_Berry is operational.

So all-in-all, not a very productive period.  Hopefully many of the problems will get fixed quickly!